Evaluate SI + y2 - zdS where S is the part of the cone 22 = x2 + y2 that lies between the planes z = 1 and 2 = 4. Preview
Answer #1

2 +y et g(x,y) = x² + y²_22 x : 2x 3 - 22 = - 22 و (2-هل : ولی او روبر (2-1) ال = هل (2) + 2 + (2-1 الاء Z + (=- ال - على رح+\\ 2 6-7) 2017 99 4 21 2) 22 (43) and e ملے - 22 ( * )( 20 (( N ۱۱ که مل 22 وهل (3)- ( do S 4 \ 2 x 2817 Cel 20 را 2 x 28/2 [